OpenStudy (xxorejina):

Which is the equation of the line that contains points (0, 5) and (5, 8)? a.) y-5=5/8(x-8) b.) y-8=5/3(x-5) c.) y-8=3/5(x-5) d.) y-5=3/5(x-8)

pooja195 (pooja195):

Start by finding the slope \[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

OpenStudy (xxorejina):

so, 3/5? @pooja195

pooja195 (pooja195):

Good now use this \[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

OpenStudy (xxorejina):

y-5=3/5(x-0) ??? @pooja195